# Towards automated formal security analysis of SAML V2.0 Web Browser SSO standard — the POST/Artifact use case

### v1

@@ VII Conclusions and future Work
We deliver a formal model of the use case from the SAML V2.0 Web Browser SSO Profile: SP-initiated SSO with POST/Artifact Bindings. While building our model we used specific formal model of TLS communication which enables Dolev-Yao adversary to exchange messages with honest agents by compromising an honest agent. We analyzed 8 protocol variants that emerged from differently defining three sensitive optional features. We tested these variants against comprehensive set of authentication, secrecy and freshness properties. Automatic verification analysis has found the artifact-session confusion attack which breakes the IdP-SP non-injective agreement property. Although by this attack the adversary could gain access to arbitrary user resource, we argued that there is no practical impact of this attack. But this attack emphasises that care must be taken while implementing SAML processing rules completely.

### v2

@@ VII Conclusions and future Work
We take a step towards comprehensive formal analysis of SAML V2.0 Web Browser SSO Profile by proposing a simple methodology and use it to formally verify an SP-initiated SSO with POST/Artifact Bindings use case.
Before building the model, we identify sensitive optional features and sensitive under-specified mechanisms and analyze eight different protocol variants to test if seemingly functional choices can impact the security of the proposed protocol. The model allows for almost trust relationships, and includes a Dolev-Yao adversary that completely controls the network and can compromise arbitrary participants.
We consider a comprehensive set of authentication, secrecy and freshness properties and, through machine-verifiable formal analysis, show that the considered use case does indeed provide the most important security properties expected of a Web SSO protocol.
We do discover minor weaknesses, including the artifact-session confusion attack which breaks the IdP-SP non-injective agreement property. We argue that this vulnerability holds no practical impact when the SP implements SAML in accordance with standard. Nevertheless, we suggest a slight modification to the standard and formally verify that it does indeed address the discovered weakness.